George is 5 times older than his son. In 15 years he will only be 2 times older than his son. How old will the son be in 15 years?
Answer by lwsshak3(11628)

George is 5 times older than his son. In 15 years he will only be 2 times older than his son. How old will the son be in 15 years?
***
let x=son' s age
5x=George's age
..
after 15 yrs
x+15=son's age
5x+15=George's age
5x+15=2(x+15)
5x+15=2x+30
3x=15
x=5
x+15=20
How old will son be in 15 years? 20